### v1

# Impacts of Dynamical Dark Energy on the Neutrino Mass Constraints

Abstract
The difference between the total neutrino mass estimates ($\sum m_{\nu}$) derived from cosmological data within the standard $\Lambda$CDM model and those obtained from terrestrial particle physics experiments underscores the need to explore alternative scenarios. Recent analyses have shown that a dynamic dark energy modeled by the CPL parameterization of the dark energy equation of state (EoS) can ease the constraints on $\sum m_{\nu}$, thus alleviating this tension. This study investigates the robustness of this discrepancy by assessing the extent to which the CPL assumption influences the results. We examine how other EoS parameterizations, such as the Barboza-Alcaniz (BA) and Jassal-Bagla-Padmanabhan (JBP) parameterizations, affect $\sum m_{\nu}$ estimates. We perform a Markov Chain Monte Carlo (MCMC) analysis combining the latest baryon acoustic oscillation data from DESI with the Planck 2018 cosmic microwave background data, which includes information on temperature, polarization, and lensing, as well as the Pantheon+ Type Ia supernovae observations. While both the BA and JBP parameterizations can also resolve the tension, our results show a correlation between the dark energy EoS and the constraints on neutrino mass.
I Introduction
In the standard cosmology, a dark energy (DE) component drives the current cosmic acceleration. This component possesses negative pressure and is characterized by an Equation of State (EoS) parameter defined as $w=p/\rho$, where $p$ is the dark energy pressure and $\rho$ its energy density. A natural theoretical representation of DE is the vacuum energy density, also known as the cosmological constant $\Lambda$, for which $w=-1$. However, despite being able to explain most of the present-day observational data, the standard $\Lambda$-Cold Dark Matter ($\Lambda$CDM) model faces tensions in measurements of some cosmological parameters. The most prominent of these discrepancies is a tension of $\sim 5\sigma$ between the Hubble Constant $H_{0}$ measurements in the late- and early-time Universe. Furthermore, $\Lambda$ is associated with well-known fundamental problems in theoretical physics [1, 2, 3, 4].
In this context, various alternative proposals have been suggested to explain the current cosmic acceleration beyond the cosmological constant (see e.g. [5, 6, 7, 8, 9, 10, 11]). In general, such proposals assume the existence of scalar fields that have not yet reached their ground state or modifications in the Friedmann equations, implying time-dependent EoS parameters $w(a)$. Phenomenologically, several authors have also proposed $w(a)$ parameterization [12, 13, 14, 15, 16], which insert the $\Lambda$CDM model in a more general framework, opening the possibility of checking deviations from the standard case.
To assess the observational viability of dynamical dark energy (DDE) models, several probes have been utilized from both the early and local Universe [17, 18, 19, 20, 21, 22, 23]. It is important to note that CMB data alone cannot constrain such models tightly. This limitation is due to a geometric degeneracy between the parameters that affect the late-time Cosmology and the angular diameter distance to the surface of the last scattering. Therefore, this degeneracy must be broken using local Universe observables to constrain this class of models. Although incorporating information such as SN data alongside CMB observations helps tighten the constraints, no preference for this class of models has been shown at first. This remained the case until very recently, when BAO observations from the Dark Energy Spectroscopic Instrument (DESI), combined with Planck CMB and Pantheon+ SN data, showed a significant preference for DDE using the Chevallier-Polarski-Linder (CPL) EoS parameterization, $w(a)=w_{0}+w_{a}(1-a)$ [24]. Even more recently, it has been shown that this preference is not limited to the CPL but extends to other parameterizations as well. This result further reinforces the preference indicated by parametric analyses of the DESI and SNe data for DDE models [25] [^1] This preference for a DDE is not confirmed by non-parametric (model-independent) analyses of the current BAO and SN data, as shown in [26, 27] (see also [28])..
Beyond the DDE discussion, the late-time universe is plagued by further uncertainties related to the background and perturbations evolution, and the neutrino physics is of particular interest in this picture. In the standard scenario, outlined by the Standard Model of fundamental particles (SM), neutrinos are massless particles. Such a conclusion is challenged by the number density measurements of the solar and atmospheric neutrinos, pointing to an oscillatory behavior of the three flavors. Oscillation experiments are insensitive to the absolute mass scale of the neutrinos, but useful to probe their squared mass differences. The latter results from atmospheric neutrinos point to $|\Delta m_{32}^{2}|\approx 2.4\times 10^{-3}\;\text{eV}^{2}$, while from solar neutrinos one obtains $|\Delta m_{21}^{2}|\approx 7.4\times 10^{-5}\;\text{eV}^{2}$ [29]. This configuration enables at least two non-relativistic species of neutrinos. Lower limits on the sum of their masses can also be estimated: $\sum m_{\nu}>0.06\;\text{eV}$ for the Normal Hierarchy schemes (NH), while the Inverted Hierarchy (IH) requires $\sum m_{\nu}>0.10\;\text{eV}$.
Back to the cosmological perspective, neutrinos have a well-marked role in the expansion of the cosmos. They act as radiation in the very early universe, later changing to a hot dark matter behavior after the non-relativistic transition. At the perturbative level, the neutrinos dispersion acts to suppress the growth of perturbation at small scales, characterized by their free-streaming and, ultimately, by the sum of their masses. One of the current tightest constraints, $\sum m_{\nu}<0.07\;\text{eV}$ at 95% confidence level (C.L), comes from the combination of CMB and DESI BAO data within the $\Lambda$CDM framework, showing a clear preference for the NO [24, 30] (see also [31]). More recently, an upper limit of $\sum m_{\nu}<0.042$ eV was obtained using CMB data combined with DESI BAO, the galaxy cluster angular diameter distance, and an $H_{0}$ prior from SH0ES [32].
These constraints are significantly relaxed within a DDE framework modeled by the CPL parameterization, yielding $\sum m_{\nu}<0.15\;\text{eV}$ at 95% C.L [24]. Therefore, the current cosmological limits for the $\Lambda$CDM are on the verge of excluding both inverted and normal ordering, thereby generating tension with the results from particle physics experiments [33, 32]. This work examines how other well-established EoS parameterizations, such as the Barboza-Alcaniz (BA) and Jassal-Bagla-Padmanabhan (JBP) parameterizations, impact the $\sum m_{\nu}$. In our analysis, we perform a Markov Chain Monte Carlo (MCMC) analysis combining the latest CMB, SN, and DESI BAO data. The paper is organized as follows. In Section 2, we detail the functional forms of the DDE models considered. In Section 3, we present the methodology adopted and the results obtained. In Section 4, we present and discuss our main conclusions.
II Dynamical Dark Energy Models
DDE models affect the evolution of the Universe in two primary ways: through the expansion rate and through the cosmological perturbations that give rise to structures. In the case of the Universe’s expansion rate, assuming a statistically homogeneous and isotropic Universe that follows the Friedmann-Lemaître-Robertson-Walker (FLRW) metric, we have:
$H^{2}=H_{0}^{2}\left[\rho_{r}a^{-4}+\rho_{m}a^{-3}+\rho_{de}a^{-3}\exp{[\Upsilon(a)]}\right]\;,$ (1)
where $w(a)$ represents the EoS as a function of the scale factor and $\Upsilon(a)=\left[\int_{1}^{a}-3\left(\frac{w(a)}{a}\right)da\right]$. For the cosmological perturbations the first and most obvious impact of a time-evolving EoS is on the density perturbation and on the fluid velocity. If we assume a Newtonian gauge,
$ds^{2}=-(1+2\Psi)dt^{2}+a^{2}(t)(1-2\Phi)\delta_{ij}dx^{i}dx^{j},$ (2)
then, the Continuity and Euler equations are written as
$\dot{\delta}+3\mathcal{H}(c_{s}^{2}-w)\delta+(1+w)\left(\theta-3\dot{\Phi}\right)=0,$ (3a)
$\dot{\theta}+\mathcal{H}(1-3w)\theta-\frac{k^{2}c_{s}^{2}\delta}{1+w}-k^{2}\Psi=0.$ (3b)
The dark energy also affects the adiabatic sound speed of the fluid $c_{s}^{2}$. To investigate the impact of of DDE models on the sum of neutrino mass, we will use two functional forms for the EoS parameter:
  • The Jassal-Bagla-Padmanabhan parameterization [14]
    $w(a)=w_{0}+w_{a}a(1-a)$ (4)
The parameterization above consists of a combination of linear and quadratic terms in the scale factor. In the current epoch, the term $-w_{a}a^{2}$ becomes comparable to $w_{a}a$, leading to differences at low-redshifts compared to the CPL model.
  • The Barboza-Alcaniz parameterization [15]
    $w(a)=w_{0}+w_{a}\frac{1-a}{a^{2}+(1-a)^{2}}$ (5)
This parameterization exhibits a linear behavior at low-redshifts. Unlike the CPL parameterization, which blows up exponentially as the scale factor approaches infinity (for $w_{a}>0$), the BA parameterization remains well-behaved at all times, while still allowing for deviations from the CPL scenario [^2] Among other well-known EoS parameterizations, the BA gives the most significant improvement over the $\Lambda$CDM model when analyzed in the context of current CMB, BAO, and SN observations [25]..
III Analysis and Results
To obtain constraints on the cosmological parameters for the DDE models above, we perform a Monte Carlo Markov Chain (MCMC) numerical analysis using the Planck 2018 temperature and polarization data (including low and high-multipole modes and lensing) [34], along with SN data from Pantheon+ [35, 36] and geometric BAO information from the DESI collaboration [24]. For this, we implement the two parameterizations in the Boltzmann solver code Cosmic Linear Anisotropy Solving System (CLASS) [37], in which we interfaced with the Cobaya code to perform the MCMC statistical analysis [38, 39].
Figure 1: Confidence contours at $68\%$ and $95\%$ C.L. for the JBP and BA parameterizations with Planck 2018 temperature, polarization and lensing together with Pantheon+ SN and DESI BAO data.
Figure 1 shows the 68% and 95% confidence level contours obtained from our analysis. It is immediately evident that the main differences between them lie in the EoS parameters. It shows significant differences, particularly in $w_{0}$, $w_{a}$, and $\sum m_{\nu}$. Table 1 presents each model’s mean values and associated errors. The $2\sigma$ limits are provided for the sum of the neutrino masses, while the $1\sigma$ limits are shown for the other parameters. For the BA model, we find an upper 95% one-tail limit of $\sum m_{\nu}<0.18$ eV. The JBP model provides a more restrictive upper limit, with $\sum m_{\nu}<0.13$ eV. Using the same data set, the CPL parameterization yields an upper bound of $\sum m_{\nu}<0.18$ eV  [^3] Our neutrino mass constraints for the CPL model are slightly different from the one reported in the DESI paper [24]. Here, we use CMB data from the Planck baseline likelihood only, while the DESI collaboration uses Planck combined with Data Release 6 of the Atacama Cosmology Telescope.. These results demonstrate that the $\sum m_{\nu}$ constraints depend on the parameterization adopted in the analysis. However, for those considered in this study, the limits obtained alleviate the tension and agree with the mass hierarchy scheme [^4] For completness, we also conduct analyses using a logarithmic parameterization defined as $w(a)=w_{0}+w_{a}\ln(a)$. Although not presented here, the results derived from this approach are similar to those obtained by the BA parameterization..
   Parameter BA JBP CPL
$h$ $0.6808\pm 0.0073$ $0.6792\pm 0.0071$ $0.6801\pm 0.0071$
$\Omega_{m}$ $0.3080\pm 0.0071$ $0.3082\pm 0.0069$ $0.3085\pm 0.0069$
$\Omega_{b}$ $0.0484\pm 0.0011$ $0.0486\pm 0.0010$ $0.0484\pm 0.0010$
$n_{s}$ $0.9657\pm 0.0040$ $0.9667\pm 0.0038$ $0.9658\pm 0.0037$
$\sigma_{8}$ $0.816\pm 0.012$ $0.814\pm 0.011$ $0.815^{+0.012}_{-0.011}$
$\sum m_{\nu}$$(\mathrm{eV})$ $0.072^{+0.11}_{-0.074}$$(2\sigma)$ $0.050^{+0.083}_{-0.051}$$(2\sigma)$ $0.070^{+0.11}_{-0.073}$$(2\sigma)$
$w_{0}$ $-0.858\pm 0.055$ $-0.795\pm 0.089$ $-0.813\pm 0.064$
$w_{a}$ $-0.36^{+0.17}_{-0.13}$ $-1.30\pm 0.60$ $-0.73^{+0.34}_{-0.26}$
[…]
In Figure 2, we show the $w_{0}$-$w_{a}$ plane. When treating neutrino mass as a free parameter, the JBP model progressively aligns with the $\Lambda$CDM scenario, in contrast to the BA model, which remains more than $2\sigma$ away. Similarly, the CPL model shows that allowing for free neutrinos has a slight influence on its analysis, consistent with the results reported by the DESI collaboration. This result suggests that relaxing constraints on the dark energy parameters may help constrain the neutrino mass, although the correlation is not particularly strong.
Figure 3 shows the EoS evolution with redshift for each model, using Planck CMB temperature, polarization, and lensing data, along with Pantheon+ SN and DESI BAO. They behave similarly at low redshifts, remaining in the quintessence region, i.e., $w(z)>-1$. At high redshifts, there are significant differences. The BA model crosses at a slightly lower redshift and remains in the phantom region near $w(z)=-1$. Interestingly, for the JBP model, it tends to cross $w(z)=-1$ twice. It leaves the quintessence region and remains very close to $w(z)=-1$ with a trend toward the phantom region until it returns to the quintessential region for high $z$ values. These EoS behaviors for these models are similar to those presented in the work [25], where a similar analysis was performed with the neutrino mass fixed. The most significant difference is in the JBP behavior, which shows a larger error in $w(z)$.
Figure 2: $w_{0}-w_{a}$ plane for the JBP, BA and CPL parameterizations with Planck 2018 temperature, polarization and lensing together with Pantheon+ SN and DESI BAO data. The dashed line intersection represents the SCM values, $w_{a}=0$ and $w_{0}=-1$. The filled contours are for the models assuming the neutrino mass as a free parameter (Blue, Red and Green), while the black empty contours (shown here for completeness) stand for the JBP model with a fixed neutrino mass.
Figure 3: EoS evolution obtained for Planck+DESI+Pantheon+ data assuming the BA (Blue) and JBP parameterizations (red). The dashed line indicates the $\Lambda$CDM model, $w(z)=-1$.
IV Final Remarks
Neutrinos are one of the most promising probes to unveil aspects of the physics at high energies. For instance, in the well-known seesaw paradigm [40, 41, 42, 43], the magnitude of the neutrino masses is dictated by a new energy scale, associated with lepton number violation. The mechanisms that lead to this symmetry breaking process may be associated with Grand Unification theories [44], and also may offer a solution to the fermion/anti-fermion unbalance of the universe [45]. Meanwhile, the micro-physical aspects of the standard (active) and sterile neutrinos are set by the Yukawa structure of the model’s Lagrangian, which may be constrained by the oscillation process observed in the atmospheric and solar neutrinos and also set the possible signatures left by new particles in related experiments [46, 47].
In this regard, the complementarity of cosmological and terrestrial observations is an important aspect for constraining the theory. While neither type of experiment can determine their fundamental properties independently, their combination can help narrow down the possible mass scale and provide insights into their interactions. Thus, as the tension between neutrino mass measurements derived from cosmological data in the $\Lambda$CDM framework and those obtained from precise ground-based experiments becomes increasingly apparent, it is crucial to explore alternative models to address this inconsistency. In this context, the DDE scenario modeled by the CPL parameterization has demonstrated the potential to resolve this discrepancy, yielding $\sum m_{\nu}<0.15\;\text{eV}$ at a 95% (CL) [24]. Additionally, parametric analyses of recent data from DESI indicate a slight preference for this model compared to the $\Lambda$CDM framework.
In this work, we investigated the robustness of the $m_{\nu}$ tension and the results obtained from the CPL model. By performing a MCMC analysis with the most recent available BAO data from DESI along with temperature, polarization, and lensing data from Planck and Type Ia SN from Pantheon+, we studied how different well-known EoS parameterizations, such as BA and JBP parameterizations, influence the estimates of $\sum m_{\nu}$. We found that the BA and JBP follow the same trend as the CPL by helping to relax the constraint on the sum of neutrino masses. The neutrino mass upper limit obtained for the BA is $\sum m_{\nu}<0.18$ eV while the JBP parameterization provides a more restrictive upper limit of $\sum m_{\nu}<0.13$ eV, a difference of $\sim 30\%$. We also note that the main difference lies primarily in the constraints on the EoS parameters. This result suggests that a larger uncertainty on $w_{0}$ and $w_{a}$ values leads to tighter constraints on the neutrino mass. For the JBP parameterization, we obtain constraints on $w_{0}$ and $w_{a}$ that are compatible with the $\Lambda$CDM model values within $2\sigma$. However, this is not the case for the BA, which remains more than $2\sigma$ away.
Finally, our results confirm the ongoing tension in the measurements of $\sum m_{\nu}$ within the context of the standard cosmological model, as reported by [33, 32]. They also demonstrate that dynamical dark energy models can effectively alleviate this issue. Furthermore, our analysis indicates that the degree to which constraints on $\sum m_{\nu}$ are relaxed depends moderately on the chosen parameterization.
Acknowledgements
GR is supported by the Coordenação de Aperfeiçoamento de Pessoal de Nível Superior (CAPES). JR thanks the Fundação de Amparo à Pesquisa do Estado do Rio de Janeiro (FAPERJ) grant No. E-26/200.513/2025. JSA is supported by CNPq grant No. 307683/2022-2 and FAPERJ grant No. 259610 (2021). We also acknowledge the use of the Cobaya, CLASS and GetDist. This work was developed thanks to the computational support of the National Observatory Data Center (CPDON).